# TOG: Training-free Task-oriented Grasp Generation

### v2

@@ Abstract
We presents a training-free pipeline for task-oriented grasp generation (TOG) that integrates pre-trained grasp generation models with vision-language models (VLMs). Beyond focusing on stability, our method incorporates task-specific requirements by leveraging the semantic reasoning capabilities of VLMs. We evaluate different strategies for integrating the VLM with a pre-trained grasp generation model, demonstrating through extensive empirical studies that this simple yet effective approach performs robustly in both simulated environments and real-world robotic experiments. Our results highlight the potential of VLMs to enhance task-oriented manipulation, offering valuable insights for future research in robotic grasping and human-robot interaction. Codes can be found at: https://github.com/StevenLiudw/TOG
@@ I Introduction
Several grasp generation methods have been introduced in recent research [2, 7, 4, 3], with the majority focusing on achieving stable, collision-free grasps. While stability and collision avoidance are critical, human grasping behavior is inherently task-driven. For humans, grasping is not just about securing an object safely but about fulfilling a specific goal. For instance, when picking up a cup to drink, a person typically grasps it by the handle, even though other stable, collision-free grasps — like holding the cup’s body — are possible. This task-oriented approach highlights that grasp selection is based on the intended use of the object — an idea not fully embraced by traditional robotic grasping techniques.
@@
Figure 1: Given a specific task, TOG can generate feasible grasps according to the task specifications by leveraging the combined capabilities of Contact Grasp Net (CGN) and a Vision-Language Model (VLM).
@@
In this project, we propose a training-free pipeline for task-oriented grasp generation (TOG) by integrating a pre-trained grasp generation model with a VLM. Our key insight is that for more capable VLMs, allowing them to freely specify grasp points and match them to generated grasps yields the best performance. In contrast, for less powerful models, a more constrained approach — where the VLM selects from pre-filtered grasps — leads to better results.
Figure 2: Overview of the system. The depth image is unprojected into a point cloud, which is processed by a grasp generation model (e.g., Contact GraspNet) to produce a set of unconditional grasps. The top K grasps are selected based on confidence and further refined by a motion planner to ensure trajectory feasibility. These filtered grasps are then evaluated using a vision-language model (VLM) to select the best grasp for the task.
We extensively evaluated and compared different approaches to integrating grasp generation models with VLMs in both a simulation environment [17] and real robot settings. Our experiments demonstrate that this simple yet effective strategy is highly robust, achieving an average 70% success rate on real robots. This study provides valuable insights into how pre-trained models and VLMs can be seamlessly integrated to enhance task-oriented grasping.
The main contributions of this work are as follows:
  • We propose a simple, flexible, and training-free pipeline for task-oriented grasp generation, along with an automated evaluation system.
  • We conduct extensive evaluations of different integration approaches, analyze failure modes in detail, and provide insights to guide future research.
@@ II-B Task-oriented Grasp Generation
In contrast, task-oriented grasp generation focuses on predicting grasps that are optimized for a specific task. Researchers in [1, 15] introduced data-driven approaches, where large datasets were collected to train neural networks capable of predicting task-oriented grasps. More recently, several studies have explored the integration of large language models (LLMs) and vision language models (VLMs) to enhance task-oriented grasp generation. For example, [13, 14] leveraged language information alone, while [12] required the training of additional grasp generation models. [10] employed differentiable classifiers to guide the grasp sampling process; however, while this approach is training-free, it is not directly applicable to more general scenarios where gradient information is not available.
@@ III-B System Overview
Given an RGB and depth image as input, our pipeline begins by processing the depth image to generate a point cloud through unprojection. This point cloud is then input into a pre-trained grasp generation model to produce a set of candidate grasps $\{G_{i}\}_{i=1}^{N}$. For our experiments, we utilized Contact GraspNet [3] due to its robustness. However, our pipeline is flexible and can accommodate other, potentially more advanced, grasp generation models.
@@ IV-A Experimental Setup
TABLE II: Three category of tasks using in simulation experiment
@@
We employed Contact GraspNet (CGN) [3] to generate candidate grasps for each object. CGN outputs grasps with associated quality scores, which were filtered using a motion planner to remove kinematically infeasible or unreachable grasps. We then followed the methodology described in Section III to generate a diverse set of grasp candidates. For the $k$-means algorithm, we set $k=3$, as it is robust across different VLMs.
@@
  • Qwen-VL [28]: An open-source VLM known for its strong language understanding and visual reasoning capabilities.
  • Molmo [29]: A VLM designed for robotic applications, particularly point-based interactions, making it well-suited for the PSI strategy.
@@ V Real-world Experiment
To test if our proposed method works in the real world, we extends the evaluation of our proposed method beyond simulation environments to assess its performance in real-world robotic manipulation tasks.
@@ V-A Experimental Setup
Our hardware configuration employed a Franka Emika Panda robotic arm equipped with an Intel RealSense L515 RGB-D camera mounted on the end-effector link (eye-in-hand). We use the RGB and depth image to generate the point clouds. We use a single singe image to query the VLM and generate the grasps, thus mimic our setup in the simulator.
We utilized Contact GraspNet (CGN) [3] to generate unconditional grasps for each object and used Gemini-2.0-Pro [33] as the VLM model. We use the CPG (Contact Point Generation) strategy as it performs the best III, in which we directly ask the VLM to generate the best grasp point that satisfying the task description. Interesting, we find that the VLM can correctly select the corresponding grasps for the target object, without using a object detection and segmentation pipeline as in the common setup, indicating the use of VLM to further simplify the grasp generation pipeline.
@@ V-B Result
Table IV presents the results of our real-world experiment. All three scenes consist of multiple objects, requiring the system to generate collision-free grasps. By integrating ContactNet’s [3] capability to generate collision-free grasps from partial point clouds with the VLM’s ability to predict grasping points based on the task description, the TOG system remains both simple and effective. Requiring only a single pair of RGB and depth images as input, it achieves an overall success rate of 70%.
@@ VI Conclusion
Despite its effectiveness, we identified several limitations, including insufficient grasp diversity, query image ambiguities, and constraints inherent to VLMs, highlighting areas for future improvement. This study underscores the potential of combining pre-trained grasp models with VLMs to enable flexible, efficient task-oriented manipulation without additional training, offering a promising framework for advancing robotic grasping capabilities.

### v3

@@ Abstract
We present a training-free pipeline for task-oriented grasp generation (TOG) that integrates pre-trained grasp generation models with vision-language models (VLMs). Beyond focusing on stability, our method incorporates task-specific requirements by leveraging the semantic reasoning capabilities of VLMs. We evaluate different strategies for integrating the VLM with a pre-trained grasp generation model, demonstrating through extensive empirical studies that this simple yet effective approach performs robustly in both simulated environments and real-world robotic experiments. Our results highlight the potential of VLMs to enhance task-oriented manipulation, offering valuable insights for future research in robotic grasping and human-robot interaction. Codes can be found at: https://github.com/StevenLiudw/TOG
@@ I Introduction
Several grasp generation methods have been introduced in recent research [1, 2, 3, 4], with the majority focusing on achieving stable, collision-free grasps. While stability and collision avoidance are critical, human grasping behavior is inherently task-driven. For humans, grasping is not just about securing an object safely but about fulfilling a specific goal. For instance, when picking up a cup to drink, a person typically grasps it by the handle, even though other stable, collision-free grasps (like holding the cup’s body) are possible. This task-oriented approach highlights that grasp selection is based on the intended use of the object.
@@
Figure 1: Given a specific task, TOG can generate feasible grasps according to the task specifications by leveraging the combined capabilities of (CGN) and a Vision-Language Model (VLM).
@@
In this paper, we describe steps towards a training-free pipeline for task-oriented grasp generation (TOG) by integrating a pre-trained grasp generation model with a VLM. Our key finding is that for more capable VLMs, allowing them to freely specify grasp points and matching them to generated grasps yields the best performance. In contrast, for less powerful models, a more constrained approach—where the VLM selects from pre-filtered grasps—leads to better results.
Figure 2: Overview of the system. The depth image is unprojected into a point cloud, which is processed by a grasp generation model (e.g., ) to produce a set of unconditional grasps. The top K grasps are selected based on confidence and further refined by a motion planner to ensure trajectory feasibility. These filtered grasps are then evaluated using a vision-language model (VLM) to select the best grasp for the task.
We evaluated and compared different approaches to integrating grasp generation models with VLMs in both a simulation environment [10] and real robot settings. Our experiments demonstrate that this simple yet effective strategy is highly robust, achieving an average 70% success rate on real robots. This study provides insights into how pre-trained models and VLMs can be seamlessly integrated to enhance task-oriented grasping.
To summarize, the main contributions of this work are as:
  • A simple, flexible, and training-free pipeline for task-oriented grasp generation;
  • An analysis of different integration approaches, including failure modes and insights to guide future research.
@@ II-B Task-oriented Grasp Generation
In contrast, task-oriented grasp generation focuses on predicting grasps that are optimized for a specific task. Recent work [5, 12] introduced data-driven approaches, where large datasets were collected to train neural networks capable of predicting task-oriented grasps. More recently, several studies have explored the integration of large language models (LLMs) and vision language models (VLMs) to enhance task-oriented grasp generation. For example, [13, 14] leveraged language information alone, while [15] required the training of additional grasp generation models. [9] employed differentiable classifiers to guide the grasp sampling process; however, while this approach is training-free, it is not directly applicable to more general scenarios where gradient information is not available.
@@ III-B System Overview
Given an RGB and depth image as input, our pipeline begins by processing the depth image to generate a point cloud through unprojection. This point cloud is then input into a pre-trained grasp generation model to produce a set of candidate grasps $\{G_{i}\}_{i=1}^{N}$. For our experiments, we utilized [4] due to its robustness. However, our pipeline is flexible and can accommodate other, potentially more advanced, grasp generation models.
@@ IV-A Experimental Setup
TABLE II: Three categories of tasks used in simulation experiment
@@
We employed (CGN) to generate candidate grasps for each object. CGN outputs grasps with associated quality scores, which were filtered using a motion planner to remove kinematically infeasible or unreachable grasps. We then followed the methodology described in Section III to generate a diverse set of grasp candidates. For the $k$-means algorithm, we set $k=3$, as it is robust across different VLMs.
@@
  • Qwen-7B [25]: An open-source VLM known for its strong language understanding and visual reasoning capabilities.
  • Molmo [26]: A VLM designed for robotic applications, particularly point-based interactions, making it well-suited for the CPSI strategy.
@@ IV-D Failure Modes
IV-E View‑Selection for Ambiguity Reduction
Figure 6: Illustration of using side cameras to obtain less ambiguous images. The first column (in red) shows the default hand-camera view, the second column (in green) shows the view with the highest confidence.
TABLE IV: Effect of view‑selection (Gemini + CPG).
Input Grasp Task Combined
Default Hand camera (orig.) 0.796 0.778 0.611
Least‑ambiguous view 0.778 0.833 0.685
As discussed in Fig. 5(b), the single frame can obscure critical affordance cues (e.g., bottle neck vs. body), leading the VLM to confuse task semantics. We hypothesize that view ambiguity comes from the image itself, not the object. If another readily available viewpoint depicts the same scene with higher semantic clarity, the VLM should perform better.
To test that, we mount four static RGB cameras around the workspace, evenly spaced at 90° on a 0.2m radius circle. For each trial we:
  1. 1.
    Capture multiple camera frames as described above
  2. 2.
    Run Detic [28] with an text prompt matching the task object to obtain the maximum class–specific confidence $\mathrm{conf}_{m}$ in each image. (Illustrated in Fig. 6).
  3. 3.
    Select the image $m^{\star}=\arg\max_{m}\mathrm{conf}_{m}$ as least ambiguous and feed it to the standard TOG pipeline.
Table IV compares results using the default hand-camera versus least-ambiguous view. While the Grasp Success Rate drops slightly, Task Compliance and Combined Success improve markedly (by 5.5 % and 7.4 %, respectively), notably improves task performance by reducing image ambiguity.
@@ V Real-world Experiment
To test if our proposed method works in the real world, we extend the evaluation of our proposed method beyond simulation environments to assess its performance in real-world robotic manipulation tasks.
@@ V-A Experimental Setup
Our hardware configuration employed a Franka Emika Panda robotic arm equipped with an Intel RealSense L515 RGB-D camera mounted on the end-effector link (eye-in-hand). We use the RGB and depth image to generate the point clouds. We use a single image to query the VLM and generate the grasps, thus mimic our setup in the simulator.
We utilized (CGN) [4] to generate unconditional grasps for each object and used Gemini-2.0-Pro [27] as the VLM model. We use the CPG (Contact Point Generation) strategy as it performs the best III, in which we directly ask the VLM to generate the best grasp point that satisfying the task description. Interesting, we find that the VLM can correctly select the corresponding grasps for the target object, without using an object detection and segmentation pipeline as in the common setup, indicating the use of VLM to further simplify the grasp generation pipeline.
@@ V-B Result
Table V presents the results of our real-world experiment. All three scenes consist of multiple objects, requiring the system to generate collision-free grasps. By integrating ContactNet’s [4] capability to generate collision-free grasps from partial point clouds with the VLM’s ability to predict grasping points based on the task description, the TOG system remains both simple and effective. Requiring only a single pair of RGB and depth images as input, it achieves an overall success rate of 70%.
@@ VI Conclusion
Despite its effectiveness, we identified several limitations, including insufficient grasp diversity, query image ambiguities, and constraints inherent to VLMs, highlighting areas for future improvement. Notably, we show that performance can be further improved through lightweight view selection to reduce perceptual ambiguity. This study underscores the potential of combining pre-trained grasp models with VLMs to enable flexible, efficient task-oriented manipulation without additional training, offering a promising framework for advancing robotic grasping capabilities.